Could you please help me with a sum that I am stuck with? Thanks in advance.
A man is 8 years older than his brother. In how many years will his age be double that of his brother?

Answer by rothauserc(4718)

let x be the age of the man's brother then the man's age is x+8
let n be the number of years
x+8+n = 2(x+n)
8 = x +n
now suppose that x = 6, then x+8 = 14
therefore n will be equal to 2 and
the man is 14+2=16 and his brother is 6+2 = 8
note that x can be any member of the set (1, 2, 3, 4, 5, 6, 7)